OpenStudy (anonymous):

Multiply. 5/4 117/20 16/9 in fraction form.

OpenStudy (anonymous):

13

OpenStudy (anonymous):

\[\frac{5}{4} *\frac{117}{20} *\frac{16}{9}=13 \]

OpenStudy (anonymous):

Thank you!